(d) How much 70% alcohol solution must be mixed with a 30% alcohol solution to produce 60 liters of a 40% alcohol solution?

Answer by checkley79(3341)

.70x+.30(60-x)=.40*60
.70x+18-.30x=24
.40x=24-18
.40x=6
x=6/.40
x=15 liters of 70% solution is used.
60-15=45 liters of 30% solution is used.
Proof:
.70*15+.30*45=.40*60
10.5+13.5=24
24=24